OpenStudy (anonymous):

sin^2x-cos^2x=0; [0,2pi)

OpenStudy (mertsj):

\[\sin ^2x-\cos ^2x=(\sin x-\cos x)(\sin x+\cos x)\]

OpenStudy (anonymous):

i dont understand…

OpenStudy (mertsj):

You don't know how to factor the difference of two squares or what?

OpenStudy (anonymous):

no i got that part, but i don't really understand whats next

OpenStudy (mertsj):

Set each factor equal to 0 and solve the resulting equation.

OpenStudy (anonymous):

so… sin^4x-2sin^2xcos^2x+cos^4x=0 then what

OpenStudy (mertsj):

\[\sin x-\cos x=0\] \[\sin x=\cos x\] \[\frac{\sin x}{\cos x}=\frac{\cos x}{\cos x}=1\] \[\tan x=1\]

OpenStudy (anonymous):

what's next after my above comment?

OpenStudy (mertsj):

I'm trying to figure out where you got that comment.

OpenStudy (anonymous):

here's what i got…sin^4x-2sin^2cos^2+cos^4x=0 1-2sin^2xcos^2x=0 sin^2xcos^2x=1/2 sin^2x=1/2 and cos^2x=1/2 (pi/6, 5pi/6) and (pi/3,5pi/3) am i doing it correctly so far?

OpenStudy (mertsj):

I can't figure out where you got that expression sin^4x-2sin^2xcox^2x+cos^4x

OpenStudy (anonymous):

i squared sin^2x-cos^2x like you said

OpenStudy (mertsj):

Where did I say to square something?

OpenStudy (anonymous):

oops, you said sin2x−cos2x=(sinx−cosx)(sinx+cosx)… so what do i do then?

OpenStudy (mertsj):

Medals 0 Set each factor equal to 0 and solve the resulting equations.

OpenStudy (anonymous):

what factors though… i don't understand where to get them

OpenStudy (mertsj):

Do you know what factors are?

OpenStudy (anonymous):

yes lol

OpenStudy (mertsj):

What are the factors of x^2-9

OpenStudy (anonymous):

3,-3

OpenStudy (mertsj):

No. The factors are (x-3)(x+3)

OpenStudy (mertsj):

If you have the equation x^2-9=0 First you factor

OpenStudy (mertsj):

And then you set each factor equal to 0 and solve each equation

OpenStudy (mertsj):

So factor sin^2x-cos^x

OpenStudy (anonymous):

i just don't see how to factor that equation

OpenStudy (anonymous):

idk if I'm having a brain fart or what but i don't know how to get them

OpenStudy (anonymous):

k well thanks for the help…..

OpenStudy (mertsj):

The same way you factored x^2-9

OpenStudy (mertsj):

The same way I showed you earlier:

OpenStudy (mertsj):

\[\sin ^2x-\cos ^2x=(\sin x-\cos x)(\sin x+\cos x)\]